OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)? y =____x + ____ What is the slope-intercept form of the function described by this table? x 1 2 3 4 y −2 −6 −10 −14 y =____x + ____ help? will fan and medal :)

OpenStudy (anonymous):

@iGreen can you help?

OpenStudy (igreen):

Okay, first we can find the slope.. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-6+2}{2-1}\) Can you solve that?

OpenStudy (anonymous):

-4\[\frac{ -4 }{ 1 }\]

OpenStudy (igreen):

Yes, now we can plug this slope and one of the points into point-slope form then simplify to get it in slope-intercept form. \(y + 2 = -4(x - 1)\) Distribute -4 into the parenthesis: \(y + 2 = -4x + 4\) Now subtract 2 to both sides, what do you get?

OpenStudy (anonymous):

y= -4x + 2?

OpenStudy (igreen):

@bcimkenz You there?